OpenStudy (anonymous):

find the HCF and LCM of 420 and 18

OpenStudy (anonymous):

you already have \[420=2^2\times 3\times 5\times 7\] \[18=2\times 3^2\] so greatest common factor is \[2\times 3=6\] and least common multiple is \[2^2\times 3^2\times 5\times 7\] whatever that is